OpenStudy (anonymous):

The square of a certain number is the same as three times the number. What is the number?

OpenStudy (anonymous):

3 or 0

OpenStudy (anonymous):

x = 0, or x = 3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw c;

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

can i get a medal

OpenStudy (anonymous):

nightmare don't harass for medals pls

OpenStudy (anonymous):

if you get one you get one if you don't, you don't.

OpenStudy (solomonzelman):

x^2=3x x^2-3x=0 x(x-3)=0 x=0 or x-3=0 x=0 or x=3 just explaining how x=0

OpenStudy (anonymous):

@plohrr ok

OpenStudy (anonymous):

THANKS!!!!!! :)

OpenStudy (solomonzelman):

Nightmare, actually, plohrr is right, you can't ask a user for medal like this, if a user wants he/she will give you the medal.

OpenStudy (anonymous):

^^

OpenStudy (solomonzelman):

\[\huge\color{blue}{ ( } \huge\color{green}{ • } \huge\color{blue}{‿ } \huge\color{green}{ • } \huge\color{blue}{ ) } \]

OpenStudy (anonymous):

wut. latex isn't working :(

OpenStudy (solomonzelman):

for you maybe not, for me it is.